OpenStudy (rane):

a golfer is teeing off on a 170m long par 3 long. the ball leaves with a velocity of 40m/s at an angle of 50 degrees to the horizontal. assuming that she hits the ball on a direct path to the hole, how far from the hole will the ball land ?

OpenStudy (rane):

@austinL

OpenStudy (austinl):

First things first. Physics problems are like essays. You need to have all your supporting information for your thesis. Meaning, you need to have everything that you know written down, draw what you have been given. And use that to form an argument as to what the answer is. Lets start with that. What do we know, and picture :)

OpenStudy (rane):

i dont understand the q' I'm confused

OpenStudy (austinl):

Lets draw a picture :) 

OpenStudy (rane):

yh?

OpenStudy (austinl):

I have to go soon, perhaps one of the other helpers here could assist? I need to prepare for class.

OpenStudy (rane):

oh ok anyone???

Parth (parthkohli):

Hi, may I help?

OpenStudy (rane):

sure

Parth (parthkohli):

Do you know a little bit of trigonometry and projectile motion?

OpenStudy (rane):

yup, i have been doing it since past 4 weeks just don't get this q;

Parth (parthkohli):

As I said in the chat, this question is just a complicated way of asking you the time taken for the projectile to end, if you read it carefully.

OpenStudy (rane):

no, its asking us the displacement between the point where it landed to from the hole

Parth (parthkohli):

whoops -- I'm too stupid :P do you know the formula for the range?

OpenStudy (anonymous):

@ParthKohli isnt it asking for the distance covered by ball??????

OpenStudy (rane):

exactly @rizwan_uet

Parth (parthkohli):

@rizwan_uet Yeah, I read it wrong. Way too stupid!

Parth (parthkohli):

Do you, anyway, know the formula for the range of any projectile? We can manage without it too, but always helps.

Parth (parthkohli):

Actually, the displacement in the horizontal direction, but you're right.

OpenStudy (rane):

\[r= u \cos \Theta * t\]

OpenStudy (rane):

i did say displacement

OpenStudy (anonymous):

formula for finding the range of projectile is R = (vi^2 sin 2theta)/g where vi = initial velocity and g= 9.8m/s

Parth (parthkohli):

Yes, but you don't need to know the time either!

Parth (parthkohli):

And it is\[R = u \sin\theta\cos\theta \times t\]

OpenStudy (anonymous):

now vi = 40m/s theta = 50 degrees and g = 9.8m/s put all these values in the above formula to get displacement @RANE

OpenStudy (rane):

i got 160.8

Parth (parthkohli):

That is correct!

OpenStudy (rane):

its not, the answer sheet says that ans = 9.38m

OpenStudy (anonymous):

what is 170m rane

Parth (parthkohli):

170 m is the distance from the hole.

Parth (parthkohli):

So subtract the range from the distance from the hole

OpenStudy (anonymous):

ok then what does par 3 long mean

OpenStudy (rane):

oh yh

Parth (parthkohli):

That's not relevant...

OpenStudy (anonymous):

ok fine

OpenStudy (anonymous):

range is 160.7 now subtract it from 170

Parth (parthkohli):

I understood the meaning of this question *now*.

OpenStudy (anonymous):

i think we should find the range

OpenStudy (rane):

ok who should i give medal to u both helped me out

Parth (parthkohli):

Anyone.

OpenStudy (rane):

don't know who,.... would u find if i give it @rizwan_uet since u r already have one?

Parth (parthkohli):

Sure.

OpenStudy (anonymous):

did you get the answer parth